solve for T please
BT^2=F^3/Z
if B=2
F=3
Z=4

Answer by checkley71(8403)

2T~2=3~3/4 OR 2T~2=27/4 OR T~2=2(27/4) OR T~2=27/2 OR T=3SQRT(3/2)